If Scores are normally distributed with a mean of 30 and a standard deviation of 5. What percent score is.



a.) Equal or greater than 30?



b.) Equal or greater than 37?



c.) From 28 to 34?

1
Expert's answer
2022-05-02T15:39:43-0400

We have a normal distribution, "\mu=30, \sigma=5."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}."


"\text{a) }z_1=\cfrac{30-30}{5}=0;
\\P(X\ge30)=P(Z\ge0)=\\
=1-P(Z<0)=\\
=1-0.5=0.5=\\
=50\%\text{ (from z-table)}."


"\text{b) }z_2=\cfrac{37-30}{5}=1.40;
\\P(X\ge37)=P(Z\ge1.40)=\\
=1-P(Z<1.40)=\\
=1-0.9192=0.0808=\\
=8.08\%\text{ (from z-table)}."


"\text{c) }z_3=\cfrac{28-30}{5}=-0.40;\\
z_4=\cfrac{34-30}{5}=0.80;\\
P(28<X<34)=\\
=P(-0.40<Z<0.80)=\\
=P(Z<0.80)-P(Z<-0.40)=\\
=0.7881-0.3446=0.4435=\\
=44.35\%\text{ (from z-table)}."